OpenStudy (jmansedg):

Find the slope of the line that contains the following pair of points. (-7, 1), (7, 8) A. 1 B. -2 C. 1/2 D. Undefined

pooja195 (pooja195):

Use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope}\]

563blackghost (563blackghost):

Use slope formula.... \(\LARGE\color{red}{\frac{ y2-y1 }{ x2-x1}}\) input the points...\(\LARGE\color{purple}{\frac{8-1}{7-(-7)}}\)

OpenStudy (jmansedg):

b? ya

pooja195 (pooja195):

Use the slope formula \[\huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope}\] \[\huge~\rm~\bf\color{blue}{ \frac{8-1=7}{7--7}=slope}\] \[\huge~\rm~\bf\color{blue}{ \frac{8-1=7}{7+7=14}=slope}\] Simplyfy the slope reduce it \[\huge~\rm~\bf\color{blue}{ \frac{7}{14}=slope}\]